x+9y =2 (1)
x= 8-9y (2)

Answer by richwmiller(17219)

x+9y =2
x= 8-9y
8-9y+9y=2
8=2
the lines do not cross
9y=2-x
y=2/9-x/9
y=-x/9+2/9
x=8-9y
9y=8-x
y=8/9-x/
y=-x/9+8/9
the lines are parallel